"\forall" x,y "\in" R, x2 +y2=9


1
Expert's answer
2022-04-25T12:32:47-0400

This statement is false. Counterexample: let "x=1, y=2". Both "x,y\in R". But, "x^2+y^2 = 1+4=5\ne9".


Answer. The statement is false.